OpenStudy (anonymous):

By trail and error with your calculator find the angle theta for which sin(theta) and theta differ by 5%. Must be done in radians.

OpenStudy (mani_jha):

Let the angle be x. \[x-\sin x=(5\sin x)/100\] \[x=105\sin x/100\] \[x=21\sin x/20\] \[\sin x=(20/21)x\] Now, differentiating this with respect to x, \[d(sinx)/dx=(20/21)(dx/dx)\] \[\cos x=20/21\] Now, find x using your calculator. Did you understand?

OpenStudy (mani_jha):

Oh! You should've posted this in the Maths group

OpenStudy (anonymous):

ok so 20/21 = 0.95 and cos 0.95=0.58 so my x= 0.58?

OpenStudy (anonymous):

or should i do cos inverse ?

OpenStudy (mani_jha):

cos inverse, of course.

OpenStudy (anonymous):

got it, thanks alot!

OpenStudy (mani_jha):

Welcome, but henceforth, don't post any maths question in physics group. By the way, I am Mani. Where are you from?

OpenStudy (anonymous):

This question is in my physics book so i thought i shud post it under physics lol Hi Mani Im Anna from california